coefficient of cubical expansion ogf a solid is (0.000027/0C). If the temperature is measured in fahreheit scale, numerical value of coefficial of linear expansion of a solid is ?

Grade:11

1 Answers

Ashrita Kanala
6 years ago
Cubical expansion of solid = linear expansion of the solid × 3--> 0.000027/°c = l.c × 3--> l.c. (alpha) = 27×10^-6 ÷ 3 = 9×10^-6Alpha in /°f = 5/9 alpha in /•c => 5/9 × (9×10^-6) => 5 ×10^-6 (or) 0.000005 /°f